Peter can buy a high-effciency furnance for $2250 that costs $412 per year to operate. He may also buy a standard furnance for $1710 that costs $466 per year to operate. Find how long it will take for the two furnances to have the same cost. If Peter intends to live in his house for 7 years which furnance should he buy? What if he wants to live in his house for 20 years?
Answer by mananth(16949)

High efficiency furnace---$2250 + $412 x : x is the number of years
Standard furnace $1710 + 466x
..
Cost equal
2250+412x= 1710+466x
2250-1710=466x-412x
540=54x
540/54 = x
x= 10 years when cost will be equal.
...
For x less than 10 years Standard furnace is economical
Plug 20 for both options
2250+412*20 =10,490 high efficiency
1710+466*20=11,030 standard
You may decide